Find the equation in the form y=mx+b
through (2,-3) and ( -3,4)
This is what i got so far :/
slope is - 7/5
y-4 = -7/5(x- -3) multiply both sides by three to get rid of it
3y-12= -4 1/5
But now I am stuck and don't know if i am doing it right

Answer by Alan3354(69443)

Find the equation in the form y=mx+b
through (2,-3) and ( -3,4)
This is what i got so far :/
slope is - 7/5
y-4 = -7/5(x- -3) multiply both sides by three to get rid of it
3y-12= -4 1/5 **** Where's the x?
=====================================
m = -7/5
y-4 = (-7/5)*(x + 3)
Multiply by 5
5y-20 = -7x - 21
7x + 5y = -1